# CMV-Fuse: Cross Modal-View Fusion of AMR, Syntax, and Knowledge Representations for Aspect Based Sentiment Analysis

### v1

@@ 1 Introduction
Despite significant advances in ABSA research, existing approaches face fundamental limitations. Context-based methods struggle to precisely associate multiple opinion terms with aspects simultaneously(e.g., associating "small" and "delicious" with "dish"). Syntactic structure-based approaches face a critical semantic-syntactic mismatch.
As observed in Figure 1, surface-level syntactic structures often fail to capture the underlying semantic relations needed for sentiment reasoning. In the dependency tree, “small” is attached to “dish” via amod, yet its position within the constituent and dependency structures still causes models to follow grammatical paths that do not reflect sentiment-bearing relationships, for example, “small” resides within the subject phrase governed by “was”, even though it semantically modifies “dish.” In contrast, the AMR graph correctly encodes “delicious-01” with dish as its ARG1, while representing “small” and “the” as semantic modifiers. However, existing knowledge-enhanced approaches struggle with inefficient integration of such structured knowledge, relying on costly subgraph sampling or weak alignment between textual and external knowledge features. As a result, they fail to efficiently leverage distinctions such as “small” indicating portion size and “delicious” indicating taste, lacking a systematic framework for combining diverse knowledge sources with linguistic representations.
To address these challenges, we propose CMV-Fuse [^1] We will release the code upon acceptance, a Cross-Modal View fusion framework that combines four complementary linguistic perspectives: Abstract Meaning Representations (AMR) for semantic abstraction, constituency parsing for hierarchical structures, dependency syntax for grammatical relationships, and semantic attention for contextual associations. Our approach emphasizes harmonizing these perspectives rather than relying on isolated linguistic views. By systematically orchestrating these representations, CMV-Fuse enhances aspect-opinion association and improves sentiment analysis in complex scenarios.
@@
    We introduce CMV-Fuse, a unified architecture that synergizes AMR semantics, constituency parsing, dependency syntax, and semantic attention through cross-view consistency constraints to bridge the semantic-syntactic gap.
@@ 2.1 Deep Learning Approaches for ABSA
Early context-based methods focused on LSTMs and attention mechanisms. tang-etal-2016-effective introduced Target-Dependent LSTM (TD-LSTM), and wang-etal-2016-attention pioneered attention-based LSTM (ATAE-LSTM). ma2017interactiveattentionnetworksaspectlevel advanced this with Interactive Attention Networks (IAN). However, these methods suffer from an aspect-opinion misalignment and limited structural awareness.
Graph-based syntactic methods emerged to capture structural relationships. zhang-etal-2019-aspect pioneered AspectGCN by applying Graph Convolutional Networks to dependency trees with aspect-focused edge pruning. sun-etal-2019-aspect extended this with Aspect-Specific GCN (ASGCN) using learnable edge weights for aspect-specific dependency adaptation. liang2022bisyn proposed the combination of constituent and dependency trees in ABSA to capture syntactic dependencies in both bottom-up and top-down directions using attention mechanisms, enabling richer contextual representations for each token. li-etal-2021-dual-graph introduced Dual-GCN with parallel syntactic and semantic graph encoders connected via BiAffine transformation. yang-etal-2020-constituency first explored constituency parsing for ABSA, demonstrating complementary phrase-level information to dependency trees. Despite advances, syntactic methods face critical semantic-syntactic mismatch where grammatical dependencies fail to align with sentiment-bearing relationships.Semantic approaches, such as Abstract Meaning Representation (AMR), provide abstraction beyond syntax. ma-etal-2023-amr introduced APARN, which improves aspect-opinion association through semantic roles. However, AMR-based approaches for ABSA remain underexplored.
@@ 2.2 Multi-View Fusion and Knowledge Integration
Recent work recognizes that single linguistic views provide incomplete information. ZHOU2020106292 pioneered knowledge graph sampling with ConceptNet subgraphs for CommonsenseGCN, while Zhong_2023 developed Knowledge Graph Augmented Network (KGAN) addressing computational complexity through efficient attention-based knowledge fusion. However, these approaches face knowledge-text alignment issues where external representations often fail to align semantically with textual features.
Multi-view fusion approaches attempt systematic integration of complementary representations. li-etal-2021-dual-graph combined syntactic and semantic graphs through BiAffine transformation, while recent work also explored contrastive learning for structural consistency. In ABSA, 10.1145/3459637.3482096 applied contrastive learning to separate sentiment features based on polarity and patterns, and LI2023110648 developed aspect-aware contrastive learning to enforce consistency between different structural views. Despite progress, existing approaches remain limited to two or three views with simple fusion strategies, lacking scalable frameworks for systematic multi-view integration.
These limitations motivate our CMV-Fuse framework, which addresses identified gaps through comprehensive AMR integration for semantic abstraction, hierarchical fusion architecture for systematic multi-view combination, and structure-aware contrastive learning for consistent cross-modal alignment, providing the first principled solution integrating AMR semantics with syntactic structures and external knowledge for robust aspect-based sentiment analysis.
@@ 3 Model Overview
As shown in Figure 2, our model consists of three key components: (1) Multi-View Graph Encoder, (2) Hierarchical Cross-Modal Fusion, and (3) Multi-view structure-aware Contrastive Alignment. The process begins with BERT contextualized representations $H^{\text{bert}}=\{h^{\text{bert}}_{1},h^{\text{bert}}_{2},\ldots,h^{\text{bert}}_{n}\}\in\mathbb{R}^{n\times d}$ from sentence-aspect pairs formatted as $x=\text{[CLS]}s\text{[SEP]}a\text{[SEP]}$, which serve as initial node features for the graph encoders that process Abstract Meaning Representations, constituency parsing, dependency syntax, and semantic attention to capture structural and contextual patterns.
@@ 3.1 Multi-View Graph Encoder Module
AMR Adjacency Matrix: Abstract Meaning Representation (AMR) provides a semantic abstraction that captures the core meaning of sentences while abstracting away surface syntactic variations. Our CMV framework incorporates AMR graphs as a complementary representation for aspect-based sentiment analysis. We utilize AMR-BART bai-etal-2022-graph to generate AMR graphs from input sentences, converting them into token-aligned adjacency matrices via concept-to-token mapping using LEAMR blodgett-schneider-2021-probabilistic.
@@
$\mathbf{A}^{dep}_{ij}=\begin{cases}1&\text{if }(w_{i},w_{j})\in\mathcal{E}_{dep}\text{ or }(w_{j},w_{i})\in\mathcal{E}_{dep}\\ 1&\text{if }i=j\\ 0&\text{otherwise}\end{cases}$ (2)
@@ 3.2.2 Graph-Specific Processing and Knowledge Integration
Each graph structure is processed independently using the unified GCN with different layer depths. The constituency representation uses $l_{c}$ layers for hierarchical parse tree processing, while the dependency representation employs $l_{d}$ layers to capture multi-hop syntactic relations. The semantic representation utilizes $l_{s}$ layers for attention-based contextual connections, and the AMR representation applies $l_{a}$ layers for semantic role and concept alignment. This layer-specific processing allows each representation to capture its unique structural patterns while maintaining consistent operations. We also integrate external knowledge represented as ${H}^{kg}=[{h}_{1}^{kg},{h}_{2}^{kg},\dots,{h}_{n}^{kg}]$.
@@ 3.3.4 Cross-Modal Global Fusion
This hierarchical fusion architecture ensures that information flows progressively from fine-grained syntactic patterns to high-level semantic understanding, while maintaining the structural inductive biases learned by each specialized GCN component.
@@ 3.4 Multi-view structure-aware Contrastive Learning
This formulation captures both global connectivity (mean attention) and peak semantic relevance (max attention), identifying tokens that serve as effective anchors for structural learning. Following the multi-head pooling strategy in MP-GCN 9691319, we select the top-$k$ tokens per sentence where $k=\max(1,\lfloor(\log_{10}(\max(2,n)))^{2}\rfloor)$. This sparse sampling approach is theoretically grounded in Bourgain’s Theorem pmlr-v97-you19b, which demonstrates that $\mathcal{O}(\log^{2}n)$ landmark points suffice to preserve essential structural properties in metric spaces.
@@ 4 Experiments
BERT devlin2019bert 85.62 78.28 77.58 72.38 75.28 74.11
dotGCN chen2022discrete 86.16 80.49 81.03 78.10 78.11 77.00
R-GAT wang-etal-2020-relational 86.60 81.35 78.21 74.07 76.15 74.88
KE-IGCN wan2023knowledge 86.70 81.05 81.06 77.89 - -
Dual GCN li-etal-2021-dual-graph 87.13 81.16 81.80 78.10 77.40 76.02
SSEGCN zhang2022ssegcn 87.31 81.09 81.01 77.96 77.40 76.02
BiSyn-GAT liang2022bisyn 87.49 81.63 82.44 79.15 77.99 76.80
CMV-Fuse (ours) 87.76 81.99 81.16 77.75 78.13 77.06
@@
4.1 Datasets and Experimental Setup
For AMR preprocessing, we use AMRBART bai-etal-2022-graph for semantic parsing and blodgett-schneider-2021-probabilistic for token alignment, with SuPar Zhang2020SuPar for dependency and constituent tree parsing.
We use the bert-base-uncased model devlin-etal-2019-bert with 768-dimensional hidden vectors and a sequence length of 100. The AMRGCN, DepGCN, ConGCN, and SemGCN layers are set to (1, 8, 6), (5, 2, 6), (3, 6, 3), and (7, 8, 6), respectively, with $\beta$ coefficients (0.5, 0.2, 0.5). Hyperparameters $\delta$ and $\gamma$ are set to 10 and 0.2. We train for 15 epochs, evaluating after each epoch. All experiments run on A100 GPUs, with each taking an average of  15 minutes. Performance is evaluated using Accuracy and macro-F1 scores.
4.2 Baseline Methods
We compare CMV-Fuse with several state-of-the-art models, including: (1) BERT devlin-etal-2019-bert, a general pre-trained model with a classification head for ABSA; (2) R-GAT wang-etal-2020-relational, which uses a relational GAT for aspect-based dependency structures; (3) DualGCN li-etal-2021-dual-graph, a dual GCN that merges features from sentences and dependency trees; (4) dotGCN chen2022discrete, which uses a latent tree structure for aspects; and (5) SSEGCN devlin-etal-2019-bert, which incorporates aspect-aware attention in GCNs.(6) KE-IGCN wan2023knowledge aims to select the highly relevant subgraphs, and proposes an interaction strategy to evaluate the interaction between external knowledge and the input text. (7) BiSyn-GAT liang2022bisyn graph-based model that captures syntactic dependencies in both bottom-up and top-down directions using attention mechanisms, enabling richer contextual representations for each token.
4.3 Main Results
Table 2 shows the experimental results on the SemEval-2014 datasets. CMV-Fuse demonstrates competitive performance across all three datasets. In Restaurant14 and Twitter, it beats all the baseline models in accuracy (87.76%, 78.13%) and achieves comparable Macro-F1 (81.99%, 77.06%). This highlights the effectiveness of our Cross Modal-View Fusion in capturing aspect-sentiment relationships.
Compared to syntax-focused models like Dual GCN (87.13%) and dotGCN (86.16%), CMV-Fuse improves by +0.63% and +1.60% respectively. Comparing it with an BiSyn-GAT, an model which also works on constituency and dependency trees, we see an accuracy improvement of +0.27% and +0.14% respectively on Restaurant14 and Twitter.
Performance variations across datasets show that: (1) Restaurant14 benefits from knowledge integration, (2) Laptop14 relies more on syntactic patterns, and (3) Twitter requires robust semantic abstraction. These findings demonstrate that CMV-Fuse effectively integrates syntactic, semantic, and knowledge views for improved aspect-based sentiment analysis.
We conduct further ablation studies (Appendix A) to assess the impact of multi-view fusion, contrastive learning objectives, and architectural choices. Results show that combining all three views (AMR, syntax, and knowledge) achieves the best performance (87.76%accuracy). Contrastive learning, especially for semantic consistency, is crucial for model performance.
@@ 5 Conclusion
This work introduced CMV-Fuse, a unified framework for aspect-based sentiment analysis (ABSA) that systematically integrates multiple linguistic perspectives: Abstract Meaning Representation (AMR), dependency and constituency syntax, semantic attention, and external knowledge. Unlike prior approaches that rely on isolated structural views, CMV-Fuse employs a hierarchical gated fusion network to harmonize local syntactic cues, intermediate semantic representations, and global knowledge signals, thereby mirroring the multi-layered nature of human language understanding. Additionally, a multi-view contrastive learning objective is proposed to align complementary structural representations and enhance cross-modal consistency without excessive computational cost. Through comprehensive experiments on benchmark datasets, CMV-Fuse demonstrates robust and consistent improvements over competitive baselines, validating the benefit of combining semantic abstraction with structural syntax and world knowledge. Our detailed ablation analyses further reveal the unique contribution of each linguistic view and confirm the framework’s adaptability to diverse sentiment contexts. Overall, this study establishes CMV-Fuse as a principled step toward more integrated, interpretable, and linguistically grounded sentiment analysis, opening avenues for future research in cross-domain generalization, multilingual adaptation, and fairness-aware model design.

### v2

@@ 1 Introduction
Context-based methods struggle to associate multiple opinion terms with an aspect simultaneously (e.g., linking “small” and “delicious” to “dish”), while syntactic structure-based approaches suffer from a critical semantic–syntactic mismatch. As shown in Figure 1, surface-level syntactic structures often fail to capture sentiment-bearing semantic relations: although “small” is attached to “dish” via the amod relation, its grammatical position can mislead models, as it resides within the subject phrase governed by “was” rather than reflecting its semantic role. In contrast, the AMR graph explicitly encodes “delicious-01” with dish as ARG1 and represents modifiers such as “small” semantically. However, existing knowledge-enhanced methods struggle to integrate such structured information efficiently, often relying on costly subgraph sampling or weak cross-view alignment, limiting their ability to distinguish sentiment attributes such as portion size and taste.
Despite recent progress, existing sentiment models lack a principled mechanism for integrating heterogeneous linguistic representations that encode complementary sentiment cues. Most prior work either optimizes a single linguistic view (e.g., syntax, semantics, or attention) or relies on ad hoc fusion strategies such as feature concatenation or task-specific heuristics, which limits cross-view alignment and generalization. Consequently, these models struggle to jointly reason over global semantic abstractions (e.g., AMR), hierarchical phrase structure (constituency), grammatical relations (dependency), and contextual associations (semantic attention), especially in complex aspect-level scenarios. To address this gap, we propose CMV-Fuse [^1] We will release the code upon acceptance, a Cross-Modal View fusion framework that independently encodes complementary linguistic views and aligns them through a hierarchical fusion architecture. CMV-Fuse adopts a modular, plug-and-play design that enables state-of-the-art single-view encoders to be systematically integrated, while explicitly modeling cross-view interactions to strengthen aspect–opinion association.
@@
    We introduce CMV-Fuse, a unified plug and play architecture that synergizes AMR semantics, constituency parsing, dependency syntax, and semantic attention through cross-view consistency constraints to bridge the semantic-syntactic gap.
@@ 2.1 Deep Learning Approaches for ABSA
Early context-based methods focused on LSTMs and attention mechanisms. Tang et al. (2016) introduced Target-Dependent LSTM (TD-LSTM), and Wang et al. (2016) pioneered attention-based LSTM (ATAE-LSTM). Ma et al. (2017) advanced this with Interactive Attention Networks (IAN). However, these methods suffer from an aspect-opinion misalignment and limited structural awareness.
Graph-based syntactic methods emerged to capture structural relationships. Zhang et al. (2019) pioneered AspectGCN by applying Graph Convolutional Networks to dependency trees with aspect-focused edge pruning. Sun et al. (2019) extended this with Aspect-Specific GCN (ASGCN) using learnable edge weights for aspect-specific dependency adaptation. Liang et al. (2022) proposed the combination of constituent and dependency trees in ABSA to capture syntactic dependencies in both bottom-up and top-down directions using attention mechanisms, enabling richer contextual representations for each token. Li et al. (2021) introduced Dual-GCN with parallel syntactic and semantic graph encoders connected via BiAffine transformation. Yang et al. (2020) first explored constituency parsing for ABSA, demonstrating complementary phrase-level information to dependency trees. Despite advances, syntactic methods face critical semantic-syntactic mismatch where grammatical dependencies fail to align with sentiment-bearing relationships. Semantic approaches, such as Abstract Meaning Representation (AMR), provide abstraction beyond syntax. Ma et al. (2023) introduced APARN, which improves aspect-opinion association through semantic roles. However, AMR-based approaches for ABSA remain underexplored.
@@ 2.2 Multi-View Fusion and Knowledge Integration
Recent work recognizes that single linguistic views provide incomplete information. Zhou et al. (2020) pioneered knowledge graph sampling with ConceptNet subgraphs for CommonsenseGCN, while Zhong et al. (2023) developed Knowledge Graph Augmented Network (KGAN) addressing computational complexity through efficient attention-based knowledge fusion. However, these approaches face knowledge-text alignment issues where external representations often fail to align semantically with textual features.
Multi-view fusion approaches attempt systematic integration of complementary representations. Li et al. (2021) combined syntactic and semantic graphs through BiAffine transformation, while recent work also explored contrastive learning for structural consistency. In ABSA, Liang et al. (2021) applied contrastive learning to separate sentiment features based on polarity and patterns, and Li et al. (2023) developed aspect-aware contrastive learning to enforce consistency between different structural views. Despite progress, existing approaches remain limited to two or three views with simple fusion strategies, lacking scalable frameworks for systematic multi-view integration
Although these prior ABSA models incorporate multiple linguistic features, their fusion mechanisms are often handcrafted or tightly coupled to specific representations, limiting extensibility and cross-view alignment. CMV-Fuse differs by introducing a modular and hierarchical fusion framework that systematically aligns heterogeneous linguistic views, allowing complementary information to be integrated without redesigning task-specific architectures.
@@ 3 Model Overview
As shown in Figure 2, our model consists of three key components: (1) Multi-View Graph Encoder, (2) Hierarchical Cross-Modal Fusion, and (3) Multi-view structure-aware Contrastive Alignment. The process begins with BERT contextualized representations $H^{\text{BERT}}=\{h^{\text{BERT}}_{1},h^{\text{BERT}}_{2},\ldots,h^{\text{BERT}}_{n}\}\in\mathbb{R}^{n\times d}$, obtained from sentence–aspect pairs formatted as $x=\text{[CLS]}\,s\,\text{[SEP]}\,a\,\text{[SEP]}$, which serve as initial node features for the graph encoders that process Abstract Meaning Representations, constituency parsing, dependency syntax, and semantic attention to capture structural and contextual patterns.
@@ 3.1 Multi-View Graph Encoder Module
AMR Adjacency Matrix: Abstract Meaning Representation (AMR) provides a semantic abstraction that captures the core meaning of sentences while abstracting away surface syntactic variations. Our CMV framework incorporates AMR graphs as a complementary representation for aspect-based sentiment analysis. We utilize AMR-BART Bai et al. (2022) to generate AMR graphs from input sentences, converting them into token-aligned adjacency matrices via concept-to-token mapping using LEAMR Blodgett and Schneider (2021).
@@
$\mathbf{A}^{dep}_{ij}=\begin{cases}1&\text{if }(w_{i},w_{j})\in\mathcal{E}_{dep}\text{ or }\\ &(w_{j},w_{i})\in\mathcal{E}_{dep}\\ 1&\text{if }i=j\\ 0&\text{otherwise}\end{cases}$ (2)
@@ 3.2.2 Graph-Specific Processing and Knowledge Integration
Each graph structure is processed independently using a unified GCN with view-specific layer depths. Specifically, the constituency, dependency, semantic, and AMR representations employ $l_{c}$, $l_{d}$, $l_{s}$, and $l_{a}$ layers, respectively, to capture hierarchical parses, multi-hop syntactic relations, contextual associations, and semantic roles. This design allows each view to model its structural characteristics while sharing a consistent encoding operation. External knowledge is incorporated as ${H}^{kg}=[{h}_{1}^{kg},{h}_{2}^{kg},\dots,{h}_{n}^{kg}]$.
@@ 3.3.4 Cross-Modal Global Fusion
This hierarchical fusion architecture ensures that information flows progressively from fine-grained syntactic patterns to high-level semantic understanding, while preserving the structural inductive biases learned by each specialized GCN component. Appendix C provides further theoretical motivation for hierarchical fusion compared to flat fusion.
@@ 3.4 Multi-view structure-aware Contrastive Learning
This formulation captures both global connectivity (mean attention) and peak semantic relevance (max attention), identifying tokens that serve as effective anchors for structural learning. Following the multi-head pooling strategy in MP-GCN Zhao et al. (2022), we select the top-$k$ tokens per sentence where $k=\max(1,\lfloor(\log_{10}(\max(2,n)))^{2}\rfloor)$. This sparse sampling approach is theoretically grounded in Bourgain’s Theorem You et al. (2019), which demonstrates that $\mathcal{O}(\log^{2}n)$ landmark points suffice to preserve essential structural properties in metric spaces.
@@ 4 Experiments
4.1 Datasets and Experimental Setup
For AMR preprocessing, we use AMRBART Bai et al. (2022) for semantic parsing and Blodgett and Schneider (2021) for token alignment, with SuPar Zhang (2020) for dependency and constituent tree parsing.
To incorporate external knowledge, we use WordNetPrinceton University (2025) as a lexical-semantic knowledge base, which is converted into relational triples encoding synonymy, antonymy, and hypernym–hyponym relations. Knowledge graph embeddings are trained on these triples using OpenKETHUNLP (2018) with DistMult, and the resulting embedding matrix serves as the external knowledge representation $h_{kg}$. Each input token is linked to its corresponding WordNet concept via lemma matching, and the associated embedding is retrieved from the trained model.
We use the bert-base-uncased model Devlin et al. (2019) with 768-dimensional hidden vectors and a sequence length of 100. For the three datasets (in the order of Table 2), the AMRGCN, DepGCN, ConGCN, and SemGCN layers are set to (1, 8, 6), (5, 2, 6), (3, 6, 3), and (7, 8, 6). Hyperparameters $\delta$ and $\gamma$ are set to 10 and 0.2. We train for 15 epochs, evaluating after each epoch. All experiments run on A100 GPUs, with each dataset taking an average of  15 minutes. Performance is evaluated using Accuracy and macro-F1 scores.
4.2 Main Results
@@
BERT Devlin et al. (2019) 85.62 78.28 77.58 72.38 75.28 74.11
dotGCN Chen et al. (2022) 86.16 80.49 81.03 78.10 78.11 77.00
R-GAT Wang et al. (2020) 86.60 81.35 78.21 74.07 76.15 74.88
KE-IGCN Wan et al. (2023) 86.70 81.05 81.06 77.89 - -
Dual GCN Li et al. (2021) 87.13 81.16 81.80 78.10 77.40 76.02
SSEGCN Zhang et al. (2022) 87.31 81.09 81.01 77.96 77.40 76.02
BiSyn-GAT Liang et al. (2022) 87.49 81.63 82.44 79.15 77.99 76.80
TextGT Yin and Zhong (2024) 87.31 82.27 81.33 78.71 77.70 76.45
S2GSL Chen et al. (2024a) 87.31 82.84 82.46 79.07 77.84 77.11
Bert+CD Tian et al. (2024) 87.32 81.94 82.25 79.65 - -
DMAN Chen et al. (2024b) 87.59 82.47 82.29 78.91 - -
MambaforGCN+BERT Lawan et al. (2025) 86.68 80.86 81.80 78.59 77.67 76.88
CMV-Fuse (ours) 87.76 81.99 82.71 79.79 78.13 77.06
@@
Table 2 shows the experimental results on the SemEval-2014 datasets. CMV-Fuse demonstrates competitive performance across all three datasets. In Restaurant14, Laptop14 and Twitter, it beats all the baseline models in accuracy (87.76%, 82.71%, 78.13%) and achieves comparable Macro-F1 (81.99%, 79.79 77.06%). This highlights the effectiveness of our Cross Modal-View Fusion in capturing aspect-sentiment relationships.
Compared to syntax-centric graph models such as Dual GCN (87.13%) and dotGCN (86.16%), CMV-Fuse improves accuracy on Restaurant14 by 0.63 percentage points and 1.60 percentage points, respectively, demonstrating the benefit of incorporating semantic abstraction and external knowledge beyond purely syntactic structures. Compared with BiSyn-GAT, which jointly models constituency and dependency trees, CMV-Fuse achieves consistent gains of 0.27 percentage points on accuracy for Restaurant14 and 0.14 percentage points for Twitter, highlighting the advantage of systematic multi-view integration over syntax-only dual-graph designs. CMV-Fuse also remains competitive with recent semantic and graph enhanced SOTA models, outperforming TextGT by 0.45 percentage points on accuracy for Restaurant14 and S2GSL by 0.40 percentage points on Twitter, while achieving the best overall performance on Laptop14 (82.71% accuracy, 79.79% Macro-F1). Overall, these results indicate that integrating AMR semantics, syntactic structure, semantic attention, and external knowledge enables more robust aspect-level sentiment reasoning across diverse datasets.
We conduct comprehensive ablation studies (Appendix B) to isolate and understand the individual contribution of each view (AMR, DEP, CON, SEM, KG), effectiveness of complementary view pairings and impact of different training objectives (e.g., alignment and contrastive losses) on performance gains. This analysis clarifies which components drive performance gains and avoids attributing improvements to the framework without evidence (results in Appendix B Tables 3 and 4). Results show that combining all three views (AMR, syntax, and knowledge) achieves the best performance (87.76% accuracy). Contrastive learning, especially for semantic consistency, is crucial for model performance.
@@ 5 Conclusion
This work presents CMV-Fuse, a unified modular ABSA framework that integrates multiple linguistic views, including AMR semantics, dependency and constituency syntax, semantic attention, and external knowledge. Unlike prior methods that rely on isolated structures, CMV-Fuse uses a hierarchical gated fusion mechanism to jointly model local syntactic cues, intermediate semantic representations, and global knowledge signals, aligning with the layered nature of language understanding. A multi-view contrastive learning objective further aligns complementary representations and improves cross-view consistency with low overhead. Experiments show that CMV-Fuse consistently outperforms strong baselines, while ablation studies confirm the complementary contributions of each linguistic view and the model’s robustness across diverse sentiment settings. Overall, CMV-Fuse advances more integrated, interpretable, and linguistically grounded sentiment analysis, with potential for cross-domain, multilingual, and fairness-aware extensions.